# Analytics of Adaptive Online Testing in Practice Over a Decade

### v1

@@ Abstract
Adaptive online testing efficiently assesses examinee proficiency by dynamically adjusting the difficulty of test items based on their performance. To achieve this, items are selected so that their difficulty closely matches the test taker’s estimated ability at each stage of the test. This alignment implies that the probability of a correct answer tends toward 0.5. However, in practical settings, this probability may not converge to 0.5 unless the test comprises a sufficiently large number of items. This could give the impression that the adaptive mechanism is not functioning properly. Nevertheless, even when the number of items is small, such as 5 or 7, the adaptive nature of the system can still be observed by examining the relationship between item difficulty and the mean estimated ability of examinees for the corresponding item.
Keywords: learning analytics, computer adaptive testing, correct answer rate, item response theory, multiple-choice problem.
@@ 1  Introduction
1.1  A brief survey on adaptive online testing
@@ 2.3  Algorithm for the adaptive online testing system
To prevent potential misunderstandings, we define two statistics concerning the estimated abilities at each problem position in the adaptive test sequence. The first is the mean of all estimates $\hat{\theta}_{j}^{(l)}$ associated with problem $j$ across test-takers and positions, while the second is the mean of the final estimate $\hat{\theta}_{j}^{(k)}$ when problem $j$ appears as the final question in a sequence. These are formally defined as follows: (1) $\hat{\mu}(\theta_{j})_{\text{all}}=E(\hat{\theta}_{j}^{(l)}|j)$, (2) $\hat{\mu}(\theta_{j})_{\text{final}}=E(\hat{\theta}_{j}^{(k)}|j)$, where $E(\cdot\mid j)$ denotes the expectation conditioned on problem $j$ being presented at a particular position in the test sequence. When asymptotic theory states that the CAR converges to 0.5 as $k\rightarrow\infty$, it implicitly assumes the use of $\hat{\mu}(\theta_{j})_{\text{final}}$. However, when the number of items administered adaptively is not sufficiently large, $\hat{\mu}(\theta_{j})_{\text{all}}$ may significantly differ from $\hat{\mu}(\theta_{j})_{\text{final}}$. In what follows, we employ $\hat{\mu}(\theta_{j})_{\text{all}}$ for direct comparison with observed CAR values.
@@ 3.2  Analytics of adaptive online testing in practice
Finally, Figure 10 presents the relationship between the mean estimated ability, $\hat{\mu}(\theta_{j})_{\text{all}}$, and the CAR value. Note that the vertical axis does not represent the mean CAR value directly, since the number of responses for each point is not considered. If we compute a weighted mean using the number of responses as weights, the resulting value matches the CAR value for each section listed in Table 4.

### v2

@@ Abstract
Adaptive online testing efficiently assesses examinee proficiency by dynamically adjusting the difficulty of test items based on their performance. To achieve this, items are selected so that their difficulty closely matches the test taker’s estimated ability at each stage of the test. This alignment implies that the probability of a correct answer tends toward 0.5. However, in practical settings, this probability may not converge to 0.5 unless the test comprises a sufficiently large number of items. This could give the impression that the adaptive mechanism is not functioning properly. Nevertheless, even when the number of items is small, such as 5 or 7, the adaptive nature of the system can still be observed by examining the relationship between item difficulty and the mean estimated ability of examinees for the corresponding item. Considering that each item typically requires about 2-5 minutes to solve, this number of items appears to be a practical choice that avoids excessive fatigue for test-takers.
Keywords: learning analytics, computer adaptive testing, correct answer rate, item response theory, multiple-choice problem, practical number of items.
@@ 1  Introduction
2  A brief survey on adaptive online testing
@@ 3.3  Algorithm for the adaptive online testing system
To prevent potential misunderstandings, we define two statistics concerning the estimated abilities at each problem position in the adaptive test sequence. The first is the mean of all estimates $\hat{\theta}_{j}^{(l)}$ associated with problem $j$ across test-takers and positions, while the second is the mean of the final estimate $\hat{\theta}_{j}^{(k)}$ when problem $j$ appears as the final question in a sequence. These are formally defined as follows:
$\hat{\mu}(\theta_{j})_{\text{all}}=E(\hat{\theta}_{j}^{(l)}\mid j),$ (3)
$\hat{\mu}(\theta_{j})_{\text{final}}=E(\hat{\theta}_{j}^{(k)}\mid j),$ (4)
where $E(\cdot\mid j)$ denotes the expectation conditioned on problem $j$ being presented at a particular position in the test sequence. When asymptotic theory states that the CAR converges to 0.5 as $k\rightarrow\infty$, it implicitly assumes the use of $\hat{\mu}(\theta_{j})_{\text{final}}$. However, when the number of items administered adaptively is not sufficiently large, $\hat{\mu}(\theta_{j})_{\text{all}}$ may significantly differ from $\hat{\mu}(\theta_{j})_{\text{final}}$. In what follows, we employ $\hat{\mu}(\theta_{j})_{\text{all}}$ for direct comparison with observed CAR values.
@@ 4.2  Analytics of adaptive online testing in practice
4.2.1   Relation between item difficulty and mean estimated ability
@@
4.2.2   Relation between the number of correct answers and the number of responses
@@
4.2.3   Relation between mean estimated ability and CAR
Next, Figure 10 presents the relationship between the mean estimated ability, $\hat{\mu}(\theta_{j})_{\text{all}}$, and the CAR value. Note that the vertical axis does not represent the mean CAR value directly, since the number of responses for each point is not considered. If we compute a weighted mean using the number of responses as weights, the resulting value matches the CAR value for each section listed in Table 4.
@@ 4.2.3   Relation between mean estimated ability and CAR
4.2.4   Relation between mean estimated ability and problem-solving time
So far, we have discussed how our proposed system maintains adaptiveness even with a relatively small number of items. It is well known that the larger the number of consecutive items, the more accurate the ability estimation tends to be. However, too many problems can cause test-takers to become fatigued and lose concentration. This suggests a trade-off between accuracy and the time required to complete the test.
Figure 11 shows the relationship between the mean estimated ability, $\hat{\mu}(\theta_{j})_{\text{all}}$, and the mean problem-solving time for each section, which supports our decision to limit the number of items: more difficult problems tend to require more time, making overly long tests impractical. The mean problem-solving time for each item, $\hat{\mu}(T_{j})_{\text{all}}$, is defined as
$\hat{\mu}(T_{j})_{\text{all}}=E(\hat{T}_{j}^{(l)}\mid j),$ (5)
where $T_{j}$ is the time duration in seconds for item $j$.
From the figure, we can see a slight positive relationship between the mean estimated ability and the solving time, although it is not particularly strong. This implies that more difficult problems generally require more time for a test-taker to tackle. For example, in section 38, the mean solving time is 301 seconds, or roughly 5 minutes. Assuming 5 such items, the total solving time would be around 25 minutes, which can be demanding in an online setting. In practice, 5 to 7 items may be a more reasonable range.
Figure 11: Relationship between mean estimated ability and problem-solving time for each section (Linear Algebra).
4.2.5   Analytics summary